# Federated Nonlinear System Identification

### v2

@@ 4.1.1 Synthetic Data
Following the approach of Xin et al. (2022), we construct heterogeneous client dynamics by perturbing a nominal system $(A_{0},B_{0})$. Particularly, for client $i$, the perturbed system matrices are as follows:
@@
4.1.2 Real-World System
We next conduct experiments across two different physical systems whose characteristics and control requirements provide contrasting evaluation scenarios. The first is a pendulum system characterized by one state and one input, targeting the estimation of two unknown parameters. Further details on the trajectory generation methodology are provided in Section 4 of Musavi et al. (2024). In all experimental settings, the control input $u_{t}$ is defined as the superposition of a deterministic policy and stochastic noise, i.e., $u_{t}=\pi(x_{t})+\eta_{t},$ where $\pi(x_{t})$ denotes the control policy evaluated at state $x_{t}$ and $\eta_{t}$ represents zero-mean i.i.d. noise. Here $\eta_{t}$ is sampled from uniform distribution distribution to model bounded stochastic disturbances.
For the pendulum, we have two nominal parameters: $A_{0}=\tfrac{1}{l}$ and $B_{0}=\tfrac{1}{ml^{2}}$. Here, $m$ denotes the mass of the pendulum and $l$ represents its length. Following the formulation of Musavi et al. (2024), we modified the trajectories by introducing the heterogeneity using the same procedure outlined in equation 9 and equation 10.
@@ 4.1.2 Real-World System
where $\alpha$ represents the angle of the rod relative to the vertical axis, $g$ is the gravity constant, $u$ is the torque input, and $w$ is the disturbance applied to this system. After discretization, the system dynamics can be rewritten in the structure of (1) with the feature vector consisting of expressions involving $\sin(\alpha)$ and $u$, all of which are analytic functions. The matrix of unknown parameters contains terms of the pendulum’s mass and the rod’s length. Each trajectory consists of $T=5$ time steps.
In the quadrotor example presented in Section 4 of Musavi et al. (2024), the control input is specified as $u_{t}=\pi(x_{t})+\eta_{t}$, where $\pi(x_{t})$ denotes the nominal controller designed according to the approach in Alaimo et al. (2013). The system comprises $13$ states and $4$ control inputs. The unknown parameter matrix $\theta_{*}$ includes $7$ parameters, such as the mass $m$ and selected elements of the inertia matrix $I$.
4.2 Estimation Error
@@
5 Results
Figure 4: Estimation error vs. global iterations for the nonlinear pendulum system using mini-batch SGD (batch size 1010). Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=2$ local updates at learning rate of $10^{-3}$. Sub-figures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=10$, $N_{i}=10$.
Figure 3 demonstrates the experimental results for a real-world, nonlinear dynamical system a pendulum. Figure 3 demonstrates the substantial impact of client participation on estimation error. With a fixed number of trajectories per client ($N_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of participating clients from 1 to 81 significantly reduces the estimation error. Figure 3 reveals the critical role of the data volume in convergence performance. When maintaining a fixed number of clients ($M_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of trajectories per client accelerates error reduction. This scaling demonstrates how federated learning enables each client to benefit indirectly from the expanded data pool through collaborative learning. Figure 3 illustrates the adverse impact of system heterogeneity on convergence behavior. Under a fixed number of clients ($M_{i}=10$) and trajectories ($N_{i}=10$), increasing the heterogeneity parameter leads to higher estimation error. As $\epsilon$ grows, it introduces greater variability into the model, which in turn makes precise estimation more difficult and increases the overall error.
Figure 4 demonstrates the system behavior under stochastic gradient descent implemented with a batch size of 10 and two local updates per client. This configuration modifies both the granularity of the local optimization and the number of iterations per aggregation round. The introduction of minibatch stochastic gradient descent (SGD), rather than full-batch updates, further injects stochasticity into the optimization trajectory and may accelerate convergence and improve generalization on unseen trajectories. Using two local training updates enhances the local update effect; this can accelerate the reduction in estimation error under heterogeneous conditions, but may also exacerbate client drift in more heterogeneous settings. We systematically analyze gradient descent and mini-batch SGD for federated nonlinear system identification, offering practical insights for choosing algorithms under communication and computational constraints. Results for the quadrotor system have been provided in the Appendix.
Figure 5 presents empirical convergence results for federated nonlinear system identification conducted on synthetic datasets, where the state-input pairs are generated from a synthetic nonlinear dynamical system. Specifically, Figure 5 illustrates that increasing the number of clients enhances estimation accuracy, given a fixed number of trajectories per client $(N_{i}=25$) and a constant heterogeneity parameter $(\epsilon=0.1)$. Similarly, Figure 5 demonstrates that augmenting the number of trajectories per client improves estimation quality when both the number of clients $(M)$ and the heterogeneity parameter $(\epsilon=0.01)$ remain fixed. Conversely, Figure 5 indicates a degradation in estimation performance as the heterogeneity parameter increases, with the number of clients $(M)$ and trajectories per client $(N_{i})$ held constant. These observations collectively underscore the trade-offs between data distribution, quantity, and heterogeneity in federated nonlinear system identification.
Figure 6 shows graphs depicting how the normalized estimation error ($e$) varies with the square root of the number of clients ($\sqrt{M}$) for different values of epsilon ($\epsilon$) and fixed number of trajectories per client $(N_{i})$ for the pendulum dynamics and synthetic data. Figure 6 adopts a configuration analogous to that in Figure 3, whereas Figure 6 employs the setup outlined in Figure 5. This empirical validation highlights how the non-asymptotic convergence rate in federated learning can be systematically improved and scaled by a factor of $\sqrt{M}$ in the low heterogeneity regime.
Figure 6 illustrates that increasing the number of local updates per communication round can reduce communication overhead by lowering the frequency of global aggregations. However, beyond a certain point, the performance gains diminish as excessive local training in each global iteration introduces significant global model divergence across clients. The experimental setup for Figure 6 is similar to that of Figure 4, except for a variation in the number of local updates.
@@ 5 Results
6 Conclusion
@@
We introduced a federated learning framework for identifying linearly parameterized nonlinear dynamical systems, with a particular focus on piecewise affine (PWA) models. Our theoretical analysis shows that the convergence error decreases as $1/\sqrt{M}$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate. We corroborate our theory on the pendulum and quadrotor nonlinear physical systems, experimentally demonstrating the improvement in convergence error at any client decreases favorably as more clients collaborate for federated system identification.
@@
Our work opens many interesting future directions. Theoretical analysis of how different optimization hyperparameters (e.g., number of local epochs in each global iteration) impact convergence is of significant interest. Adaptive batch size strategies could also balance the trade-off between gradient noise and convergence speed. A key fundamental open problem is where $\phi$ is not known a priori. One could explore scenarios where $\phi$ itself is learned collaboratively alongside the model, forming an end-to-end variant of our proposed framework. Finally, while this work focused on PWA models, a natural extension is to consider linearly parameterized nonlinear systems via Koopman theory (Brunton et al., 2022), which is based on lifting the states into an infinite-dimensional feature space, where the dynamics evolve linearly. While exact infinite-dimensional embeddings are impractical, one can explore learning a finite set of invariant eigenfunctions to approximate this transformation.
Appendix A Appendix
In this section, we provide definitions for the $\mathcal{K}$ and $\mathcal{KL}$ functions, proofs for Lemma 1, Proposition 1, Proposition 2, and Theorem 1. In the end, we provide the results for quadrotor experiments.
Definition 3 (Classes $\mathcal{K}$, and $\mathcal{KL}$).
The following function class definitions are cited from Kellett (2014).
  1. 1.
    A function $\alpha:[0,a)\to[0,\infty)$ (with $a\in(0,\infty]$) is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\alpha(0)=0$.
  2. 2.
    A function $\beta:[0,a)\times[0,\infty)\to[0,\infty)$ is of class $\mathcal{KL}$ if, for each fixed $t\geq 0$, the map $r\mapsto\beta(r,t)$ belongs to $\mathcal{K}$, and for each fixed $r\in[0,a)$, the map $t\mapsto\beta(r,t)$ is continuous, nonincreasing, and $\lim_{t\to\infty}\beta(r,t)=0$.
A.1 Proof of Lemma 1
Proof.
See Theorem 1 in Musavi et al. (2024). This lemma simply changes notation to fit the federated setting. ∎
A.2 Proof of Proposition 1
Proof.
Fix a client $i$ and a unit vector $v\in\mathbb{S}^{n_{\varphi}-1}$, and let $Y_{t}:=v^{\top}\varphi(x^{(i)}_{t},u^{(i)}_{t})$ for $t=0,\ldots,T_{i}-1$. By Lemma 1 the sequence $\{Y_{t}\}$ satisfies the $(1,s_{\varphi},p_{\varphi})$ block–martingale small-ball condition, so Proposition 2.5 of Simchowitz et al. (2018), with $k=1$ and $\alpha=1$, gives
$\Pr\!\bigl{[}\textstyle\sum_{t}Y_{t}^{2}\leq s_{\varphi}^{2}\!\bigl{(}N_{i}p_{\varphi}-2\log(1/\delta)\bigr{)}\bigr{]}\,\leq\,\delta.$
Take a $1/4$-net $\mathcal{N}$ of the unit sphere and apply the same bound to each $v\in\mathcal{N}$ with failure level $\delta/|\mathcal{N}|:=\delta$; a union bound gives that, with probability $\geq 1-\delta$, the inequality holds for every $v\in\mathcal{N}$. Because any unit $u$ is within $1/4$ of some $v\in\mathcal{N}$
$\lambda_{\min}\bigl{(}\Phi^{(i)}\Phi^{(i)\!\top}\bigr{)}\;\geq\;\frac{1}{2}\,s_{\varphi}^{2}\bigl{(}N_{i}Tp_{\varphi}-2\log(9^{n_{\varphi}}/\delta)\bigr{)}.$
If $N_{i}T\geq\tfrac{4}{p_{\varphi}}\bigl{[}n_{\varphi}\log 9+\log(1/\delta)\bigr{]}$, the component inside the bracket is at least $\tfrac{1}{2}N_{i}Tp_{\varphi}$, and replacing $p_{\varphi}$ by $1$ loses only another factor $2$, giving $\lambda_{\min}(\Phi^{(i)}\Phi^{(i)\!\top})\geq\tfrac{1}{2}s_{\varphi}^{2}N_{i}T$ with probability at least $1-\delta$. Finally, substituting $\delta/M$ for $\delta$ and union-bounding over the $M$ clients yields the solution. ∎
A.3 Proof of Proposition 2
Proof.
Conditioned on $\Phi^{(i)}$, the columns $w^{(i)}_{t}$ are independent, mean-zero, $\sigma_{w}$-sub-Gaussian vectors. For any $t=0,\dots,T-1$, the matrix-form Bernstein inequality gives
$\Pr\!\Bigl{(}\bigl{\|}W^{(i)}(\Phi^{(i)})^{T}\bigr{\|}_{2}>t\Bigr{)}\;\leq\;2\exp\!\Bigl{(}-\frac{t^{2}}{2\sigma_{w}^{2}\,\|\Phi^{(i)}\|_{F}^{2}}\Bigr{)}.$ (13)
Since $\|\phi(x,u)\|_{2}^{2}\leq b_{\phi}$, we have $\|\Phi^{(i)}\|_{F}^{2}\leq N_{i}T\,b_{\phi}\leq N_{i}T(n_{x}+n_{\phi})$. Choosing
$t\;=\;4\sigma_{w}\sqrt{\,N_{i}T\bigl{(}n_{x}+n_{\phi}+\log(2M/\delta)\bigr{)}}$
makes the right side of (13) less than or equal to $\delta/(2M)$. A union bound over $i\in[M]$ and some algebra completes the proof. ∎
A.4 Proof of Theorem 1
Proof.
Consider the relation for client $i$
$X^{(i)}_{+}\;=\;\theta^{(i)*}\,\Phi^{(i)}+W^{(i)}.$
Subtracting $\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}$ and regrouping gives
$X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}=(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}+W^{(i)}.$
Right-multiplying by $(\Phi^{(i)})^{\top}$ and summing over $i$,
$\sum_{i=1}^{M}(X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)})(\Phi^{(i)})^{\top}=\sum_{i=1}^{M}(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}(\Phi^{(i)})^{\top}+\sum_{i=1}^{M}W^{(i)}(\Phi^{(i)})^{\top}.$
Defining [^2] Because $\bar{\theta}_{\mathrm{LSE}}$ minimizes the pooled least-squares objective, the left side sum vanishes. $G:=\sum_{i}\Phi^{(i)}(\Phi^{(i)})^{\top}$ and $\bar{\theta}^{*}:=\tfrac{1}{M}\sum_{i}\theta^{(i)*},$ we obtain
$G\,(\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*})^{\top}=-\sum_{i}(\theta^{(i)*}-\bar{\theta}^{*})\Phi^{(i)}(\Phi^{(i)})^{\top}-\sum_{i}W^{(i)}(\Phi^{(i)})^{\top}.$
We can bound each term as follows:
$G\succeq\tfrac{1}{2}\,s_{\phi}^{2}N_{\mathrm{tot}}\,I$
by Proposition 1,
$\sum_{i}^{M}W^{(i)}(\Phi^{(i)})^{T}\leq 4\sigma_{w}\sqrt{N_{i}T(n_{x}+n_{\phi}+log(2M/\delta))}$
by Proposition 2, and
[…]
using Assumption 5  [^3] Notice that Assumption 5 implies that $\theta^{(i)*}-\bar{\theta}^{(i)*}||\leq\epsilon$, $\forall i\in[M]$. and that $\Phi^{(i)}(\Phi^{(i)})^{\top}||_{2}\leq N_{i}b_{\phi}$.
By multiplying both sides by $G^{-1}$, applying the bounds, taking the norm of each side gives with probability greater than or equal to $1-3\delta$,
$\bar{\theta}_{LSE}-\bar{\theta}^{*}||_{2}\leq\frac{8\sigma_{w}}{s_{\phi}^{2}}\sqrt{\frac{n_{x}+n_{\phi}+log(2M/\delta)}{\sum_{i=1}^{M}N_{i}T}}+\frac{b_{\phi}}{s_{\phi}^{2}}\epsilon.$
Since Assumption 5 implies $\|\theta^{(i)*}-\bar{\theta}^{*}\|_{2}\leq\varepsilon/2$ for $i\in[M]$,
$\|\bar{\theta}_{\mathrm{LSE}}-\theta^{(i)*}\|_{2}\leq\|\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*}\|_{2}+\varepsilon/2.$
This yields the expected bound. ∎
Figure 7: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a quadrotor using gradient descent (GD). Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$),with each client performing $K_{i}=5$ local updates at a learning rate of $10^{-1}$. Sub-figures: (a) $N_{i}=10$, $\epsilon=2$; (b) $M=5$, $\epsilon=5$.
A.5 Results for quadrotor
Another real-world system considered is the quadrotor, adapted from the model proposed in Musavi et al. (2024); Alaimo et al. (2013), with necessary modifications as described in the following. Let $p\in\mathbb{R}^{3}$ and $v\in\mathbb{R}^{3}$ denote the position and velocity of the quadrotor’s center of mass in the inertial frame, respectively. Let $\omega\in\mathbb{R}^{3}$ represent the angular velocity in the body-fixed frame, and $q\in\mathbb{R}^{4}$ represent the orientation of the quadrotor using a unit quaternion. The corresponding equations of motion for the quadrotor are given by:
$\dfrac{d}{dt}\begin{pmatrix}p\\ v\\ q\\ \omega\end{pmatrix}=\begin{pmatrix}v\\ -ge_{z}+\frac{1}{m}Qf_{u}\\ \frac{1}{2}\Omega q\\ I^{-1}(\tau_{u}-\omega\times I\omega)\end{pmatrix}+w,$
where $g$ is the gravitational constant, $m$ is the total mass of the quadrotor, $I=\emph{diag}(I_{xx},I_{yy},I_{zz})$ denotes the inertia matrix in the body-fixed frame, $f_{u}\in\mathbb{R}$ is the total thrust, $\tau_{u}\in\mathbb{R}^{3}$ is the total moment expressed in the body-fixed frame, and $e_{z}=(0,0,1)^{\top}$ is the unit vector along the inertial $z$-axis. There are seven unknown parameters in total, with heterogeneity introduced in a subset of these parameters.
$\theta_{1}$ $=\frac{1}{m}+\gamma_{1}^{(i)},\quad\theta_{2}=\frac{1}{I_{xx}},\quad\theta_{3}=\frac{I_{yy}-I_{zz}}{I_{xx}},$
$\theta_{4}$ $=\frac{1}{I_{yy}},\quad\theta_{5}=\frac{I_{zz}-I_{xx}}{I_{yy}},\quad\theta_{6}=\frac{1}{I_{zz}},$
$\theta_{7}$ $=\frac{I_{xx}-I_{zz}}{I_{zz}}.$
Here $\gamma_{1}^{(i)}\sim\mathrm{U}(0,\epsilon)$.
Figure 7 demonstrates the experimental results for quadrotor. Figure 7(a) demonstrates the impact of client participation on estimation error. With a fixed trajectory count per client ($N_{i}=10$) and system heterogeneity ($\epsilon=2$), increasing the number of participating clients reduces the error. Figure 7(b) underscores the role of data volume in convergence. Keeping the number of clients fixed ($M_{i}=5$) and heterogeneity ($\epsilon=5$) as fixed, increasing trajectories per client significantly accelerates error reduction. The length of each trajectory is set to 10 timesteps. These results follow similar trends as seen in the experiments for pendulum and synthetic datasets presented in the main paper.

### v3

@@ IV-A1 Synthetic Data
Following the approach of [21], we construct heterogeneous client dynamics by perturbing a nominal system $(A_{0},B_{0})$. Particularly, for client $i$, the perturbed system matrices are as follows:
@@
IV-A2 Real-World System
We conduct experiments across two different physical systems. The pendulum system is characterized by one state and one input, targeting the estimation of two unknown parameters [^3] Details on the trajectory generation method are in [10, Section 4]. In all both experiments, the control input is defined as $u_{t}=\pi(x_{t})+\eta_{t},$ where $\pi(x_{t})$ denotes the control policy evaluated at state $x_{t}$ and $\eta_{t}$ represents zero-mean i.i.d. noise, sampled. Here $\eta_{t}$ is sampled from a uniform distribution.
We have two nominal parameters: $A_{0}=\tfrac{1}{l}$ and $B_{0}=\tfrac{1}{ml^{2}}$ ($m$ denotes the mass of the pendulum and $l$ represents its length). Following the formulation of [10], we modified the trajectories by introducing the heterogeneity using the same procedure outlined in equations (9) and (10).
@@ IV-A2 Real-World System
After discretization, the system dynamics can be rewritten in the structure of (1).
In the quadrotor example presented in [10, Section 4], the control input is specified as $u_{t}=\pi(x_{t})+\eta_{t}$, where $\pi(x_{t})$ denotes the nominal controller designed according to the approach in [22]. The system comprises 1313 states and $4$ control inputs. The unknown parameter matrix $\theta^{*}$ includes 77 parameters to learn.
IV-B Estimation Error
@@
V Results
Figure 2 demonstrates the experimental results the pendulum. In Figure 2, as the number of clients increases, the estimation error decays faster. Figure 2 shows that data volumme per client per round slighly improves the convergence rate. Figure 2 illustrates the adverse impact of system heterogeneity on convergence behavior. Under a fixed number of clients ($M_{i}=10$) and trajectories ($N_{i}=10$), increasing the heterogeneity parameter the estimation error. Additional results for the quadrotor system can be found in the Appendix.
Figure 3 presents empirical convergence results for on synthetic datasets. Figure 3 illustrates that increasing the number of clients enhances estimation accuracy, Figure 3 demonstrates that increasing the number of trajectories per client improves estimation quality. Conversely, Figure 3 indicates a degradation in estimation performance as the heterogeneity parameter increases.
Figure 4 shows how the normalized estimation error ($e$) varies with the square root of the number of clients ($\sqrt{M}$) for the pendulum dynamics and synthetic data. This result validates that the non-asymptotic convergence rate proposed is empirically consistent in the low heterogeneity regime.
Figure 4 illustrates that increasing the number of local updates per communication round can reduce communication overhead by lowering the frequency of global aggregations. However, beyond a certain point, the performance gains diminish as excessive local training in each global iteration introduces significant global model divergence across clients.
VI Conclusion
@@
We introduced a federated learning framework for identifying linearly-parameterized nonlinear dynamical systems, with a particular focus on piecewise affine (PWA) models. Our theoretical analysis shows that the convergence error decreases as $1/\sqrt{M}$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate. We corroborate our theory on the pendulum and quadrotor nonlinear physical systems, experimentally demonstrating the improvement in convergence error at any client decreases favorably as more clients collaborate for federated system identification.
@@
Our work opens many interesting future directions. Theoretical analysis of how different optimization hyperparameters (e.g., number of local epochs in each global iteration) impact convergence is of significant interest. Adaptive batch size strategies could also balance the trade-off between gradient noise and convergence speed. A key fundamental open problem is when $\phi$ is not known a priori. One could explore scenarios where $\phi$ itself is learned collaboratively alongside the model, forming an end-to-end variant of our proposed framework. Finally, while this work focused on PWA models, a natural extension is to consider linearly-parameterized nonlinear systems via Koopman theory [23], which is based on lifting the states into an infinite-dimensional feature space, where the dynamics evolve linearly. While exact infinite-dimensional embeddings are impractical, one can explore learning a finite set of invariant eigenfunctions to approximate this transformation. A formal proof establishing that Algorithm 1 converges to the quantity characterized in Theorem 1 remains an important direction for future work. While empirical results demonstrate consistent convergence behavior, deriving rigorous theoretical guarantees is ongoing. In particular, future work will focus on establishing convergence guarantees for both GD and SGD settings, including sufficient conditions and convergence rates.
Acknowledgment
This work is supported in part by ANRF under PM ECRG Grant (ANRF/ECRG/2024/005912/ENS), in part by Wadhwani Impact Trust under Exploratory Research Grant (SB25261132EEWDWI009167), and in part by IIT Madras under New Faculty Initiation Grant (RF24251118EENFIG009167). The authors thank Nandan Sudarsanam for fruitful discussions.